# Higher Gaussian maps on the hyperelliptic locus and second fundamental form

### v2

@@ 6. Isotropic subspaces
for every $Q\in Ker(\mu_{2k})$, where $a_{ij}=-b_{g-j,g-i}$ by Remark 4.2. Recall that $Q\in Ker(\mu_{2k})$ if and only if the coefficients $a_{ij}$ satisfy the equations given in Lemma 3.7. Moreover observe that from Theorem 3.1 it follows that we have $k$ linearly independent equations in the variables $a_{u,2k+3-u}$ with $u=1,...,k+1$. We have shown that each coefficient $\lambda_{k,u}$ in equation (55) is different from zero and up to scalar it is the product $\omega_{g-3-2k+u}^{(4k+4-2u)}(p)\omega_{g-u}^{(2u-2)}(p)$. So we can multiply the sections $\omega_{g-u}$ by a non zero scalar multiple, hence equation (55) cannot be linearly dependent from the $k$ equations giving $Ker(\mu_{2k})$. Hence there exists a quadric $Q\in ker(\mu_{2k})$ such that $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})\neq 0$.
@@
The proof is very similar to the one given in Theorem 6.7, where by Remark 6.1 we must now compute
@@
By looking at the term inside the square brackets, we see that we must compute the derivatives of order $(h,l)$ of $Q$ where $h+l=4k+6$ and $20\leq l\leq 2k+$ 2. Hence we have to compute
@@
Then one immediately sees that in (56) we only have to consider the case $n+m=4k+4$. Now, after some computations which are analogous to the ones in Theorem 6.7, we get
@@
where $\alpha_{k,u}\neq 0$ for every $u$. Then we conclude as in Theorem 6.7 that
@@
Notice that saying that a nonzero element $\zeta\in H^{1}(T_{C})^{+}$ is an asymptotic direction means that the point $[\zeta]\in{\mathbb{P}}H^{1}(T_{C})^{+}$ is in the base locus of the linear space of quadrics $\rho_{HE}(I_{2})\subset Sym^{2}(H^{1}(T_{C})^{+})^{\vee}$.
@@
Let $v\in\langle\xi_{p}^{1},\xi_{p}^{3},...,\xi_{p}^{g-1}\rangle$, then $v=\sum\limits_{i=0}^{\frac{g-2}{2}}\lambda_{2i+1}\xi_{p}^{2i+1}$ and let $l$ be the maximum odd integer such $\lambda_{l}\neq 0$. We may assume $l\geq 3$, since $\rho(Q)(\xi_{p}\odot\xi_{p})=-2\pi i\mu_{2}(Q)(p)=0$ for all $Q\in I_{2}$. Then we have $v\in\langle\xi_{p}^{1},...,\xi_{p}^{l}\rangle$. Since $l$ is odd, we have $l=2k+1$ for some integer $1\leq k\leq\frac{g-2}{2}$. From Theorem 6.7 we know that there exists $Q\in Ker(\mu_{2k-2})$ such that $\rho(Q)(\xi_{p}^{2k+1}\odot\xi_{p}^{2k-1})\neq 0$ and $\rho(Q)(\xi_{p}^{i}\odot\xi_{p}^{j})=0$ for every $i,j\leq 2k-1$, $i,j$ odd integers. Then we have
@@
$+\lambda_{2k+1}^{2}\rho(Q)(\xi_{p}^{2k+1}\odot\xi_{p}^{2k+1}).$
First, notice that for every $k\leq\frac{g-2}{2}$ we have $A_{k-1,0}\subsetneq Ker_{\mu_{2k-2}}$ and $A_{k-1,0,0}\subsetneq A_{k-1,0}$. In fact, dim $Ker(\mu_{2k-2})\geq 3$, since dim $Ker(\mu_{g-4})=3$ by Theorem 3.1 and $42k-2\leq g-$ 4.
@@
$+\lambda_{2k+1}^{2}[\rho(Q)(\xi_{p}^{2k+1}\odot\xi_{p}^{2k+1})+a\rho(Q^{\prime})(\xi_{p}^{2k+1}\odot\xi_{p}^{2k+1})].$
@@
Set $W:=T_{j(C)}Y\subset H_{1}(T_{C})^{+}$ and $V:=V_{\frac{g}{2}-1}=\langle\xi_{p}^{1},\xi_{p}^{3},...,\xi_{p}^{g-1}\rangle$. We have dim$(V)=\frac{g}{2}$. We claim that $\dim(V\cap W)\leq 1$. Indeed, take $v\in V\cap W$, then $v$ is asymptotic, since $Y$ is totally geodesic. So, by Theorem 6.14 we know that $v$ is a multiple of $\xi^{1}_{p}$. Hence we have

### v3

@@ 6. Isotropic subspaces
for every $Q\in Ker(\mu_{2k})$, where $a_{ij}=-b_{g-j,g-i}$ by Remark 4.2. Recall that $Q\in Ker(\mu_{2k})$ if and only if the coefficients $a_{ij}$ satisfy the equations given in Lemma 3.7. Moreover observe that from Theorem 3.1 it follows that we have $k$ linearly independent equations in the variables $a_{u,2k+3-u}$ with u=1,,k+1u=1,...,k+1. We have shown that each coefficient $\lambda_{k,u}$ in equation (49) is different from zero and up to scalar it is the product $\omega_{g-3-2k+u}^{(4k+4-2u)}(p)\omega_{g-u}^{(2u-2)}(p)$. So we can multiply the sections $\omega_{g-u}$ by a non zero scalar multiple, hence equation (49) cannot be linearly dependent from the $k$ equations giving $Ker(\mu_{2k})$. Hence there exists a quadric $Q\in ker(\mu_{2k})$ such that $\rho(Q)(\xi_{p}^{2k+3}\odot\xi_{p}^{2k+1})\neq 0$.
@@
The proof is very similar to the one given in Theorem˜6.7, where by Remark 6.1 we must now compute
@@
By looking at the term inside the square brackets, we see that we must compute the derivatives of order $(h,l)$ of $Q$ where $h+l=4k+6$ and $0\leq l\leq 2k+2$. Hence we have to compute
@@
Then one immediately sees that in (50) we only have to consider the case $n+m=4k+4$. Now, after some computations which are analogous to the ones in Theorem˜6.7, we get
@@
where $\alpha_{k,u}\neq 0$ for every $u$. Then we conclude as in Theorem˜6.7 that
@@
Notice that saying that a nonzero element $\zeta\in H^{1}(T_{C})^{+}$ is an asymptotic direction means that the point $\zeta]\in{\mathbb{P}}H^{1}(T_{C})^{+}$ is in the base locus of the linear space of quadrics $\rho_{HE}(I_{2})\subset Sym^{2}(H^{1}(T_{C})^{+})^{\vee}$.
@@
Let $v\in\langle\xi_{p}^{1},\xi_{p}^{3},...,\xi_{p}^{g-1}\rangle$, then $v=\sum\limits_{i=0}^{\frac{g-2}{2}}\lambda_{2i+1}\xi_{p}^{2i+1}$ and let $l$ be the maximum odd integer such $\lambda_{l}\neq 0$. We may assume $l\geq 3$, since $\rho(Q)(\xi_{p}\odot\xi_{p})=-2\pi i\mu_{2}(Q)(p)=0$ for all $Q\in I_{2}$. Then we have $v\in\langle\xi_{p}^{1},...,\xi_{p}^{l}\rangle$. Since $l$ is odd, we have $l=2k+1$ for some integer $1\leq k\leq\frac{g-2}{2}$. From Theorem˜6.7 we know that there exists $Q\in Ker(\mu_{2k-2})$ such that $\rho(Q)(\xi_{p}^{2k+1}\odot\xi_{p}^{2k-1})\neq 0$ and $\rho(Q)(\xi_{p}^{i}\odot\xi_{p}^{j})=0$ for every $\leq 2k-1, i,ji,j$ odd integers. Then we have
@@
$\lambda_{2k+1}^{2}\rho(Q)(\xi_{p}^{2k+1}\odot\xi_{p}^{2k+1}).$
First, notice that for every $k\leq\frac{g-2}{2}$ we have $A_{k-1,0}\subsetneq Ker_{\mu_{2k-2}}$ and $A_{k-1,0,0}\subsetneq A_{k-1,0}$. In fact, dim $Ker(\mu_{2k-2})\geq 3$, since dim $Ker(\mu_{g-4})=3$ by Theorem 3.1 and $2k-2\leq g-4$.
@@
$\lambda_{2k+1}^{2}[\rho(Q)(\xi_{p}^{2k+1}\odot\xi_{p}^{2k+1})+a\rho(Q^{\prime})(\xi_{p}^{2k+1}\odot\xi_{p}^{2k+1})].$